OpenStudy (anonymous):

find the value of k that makes x^2 + (k+7)x+(7k+1) a perfect square trinomial

myininaya (myininaya):

If we had \[x^2+bx+c\] and we wanted to find c in terms of b so that \[x^2+bx+c\] is a complete square we would need \[c=(\frac{b}{2})^2\] So we have that: \[x^2+bx+c=x^2+bx+c=(x+\frac{b}{2})^2 \text{ since } c=(\frac{b}{2})^2 \] ---------------------------- So for you problem what should 7k+1 be in order for \[x^2+(k+7)x+(7k+1)\] for this to be a complete square?

myininaya (myininaya):

what is b here and what is c here based on what I have above?

myininaya (myininaya):

@alexeis_nicole are you there?

myininaya (myininaya):

well once you determine what b and c are just plug them into \[c=(\frac{b}{2})^2\] and solve for k

OpenStudy (anonymous):

@myininaya sorry. b = k+7 c = 7k+1 OH ! okay i get it now :)

myininaya (myininaya):

Great \[7k+1=(\frac{k+7}{2})^2\] So you know to solve this for k right?